# Runtime Stealthy Perception Attacks against DNN-based Adaptive Cruise Control Systems

### v3

@@ IV Safety Intervention Simulation
To evaluate the safety of DNN-based ACC systems under attacks, we enhance the default OpenPilot and CARLA simulation platform (see Section II-C) to be more representative of real-world ADAS, by developing a safety intervention simulator and mechanisms for priority-based dispatching of control commands to CARLA and fusion of camera and radar data (see Appendix A). An overview of the simulation platform is shown in Fig. 6 (with the orange parts representing our new implementations) and presented next.
Figure 6: Simulation platform: OpenPilot integrated with CARLA simulator and safety interventions, including driver reaction simulator, virtual Panda safety constraints checking, FCW, and AEB.
@@
To fill the gap in considering safety interventions and address the challenge of ensuring the combination of human driver and vehicle safe (see Section II-B), we implement and integrate three levels of safety interventions in the OpenPilot software (see Fig. 6), including ADAS safety features (e.g., AEB and FCW), basic car safety constraint checking on control commands, and driver interventions.
@@
AEBS (FCW and AEB) Simulator. To design and test the AEBS mechanisms in simulation, we thoroughly review the regulations and requirements concerning AEBS [78] [14] [5] and adhere to UN Regulation No. 152 [14]. We adopt and implement a time-to-collision (TTC) based phase-controlled AEBS [20] in our platform.
@@
The AEBS processes inputs from LVD outputs (after sensor fusion), including relative distance ($RD$) and relative speed ($RS$), and current speed $V_{Ego}$ (see Fig. 1). The average driver reaction time ($T_{react}$) is set to the commonly used constant value of 2.5s [97, 76]. Time thresholds, namely $ttc$ (time to collision), $t_{fcw}$ (FCW time), $t_{pb1}$ (1st phase partial brake time), $t_{pb2}$ (second phase partial brake time), and $t_{fb}$ (full brake time), are then calculated as follows:
$ttc$ $=RD/RS$ (12)
$t_{fcw}$ $=T_{react}+V_{Ego}/4.5$ (13)
$t_{pb1}$ $=V_{Ego}/2.8;t_{pb2}=V_{Ego}/5.8;t_{fb}=V_{Ego}/9.8$ (14)
@@
As shown int Fig. 8, when $ttc$ falls below $t_{fcw}$, $t_{pb1}$, $t_{pb1}$, and $t_{pb1}$, a corresponding action (warning or brake with 90%, 95%, 100% force) is executed. Applying the brake value blocks other controls from the ADAS. More details of AEBS design and testing are provided in Appendix IX.
@@
In reality, when OpenPilot is installed on a car model, some models lose the AEBS functionality [9], while others retain it. Also, AEBS might rely on a separate ADAS camera [82], distinct from the OpenPilot camera, and a potential compromise of the AEBS camera data is possible. Thus, we consider three scenarios for AEBS interventions: (1) AEBS enabled, and AEBS camera data is uncompromised; (2) AEBS enabled, but AEBS camera data is compromised; and (3) AEBS is disabled (see Section V-C2 and Table VII).
@@
Safety Constraint Checker. The OpenPilot safety mechanisms are implemented in its control software and the Panda CAN interface. Panda is a universal OBD adapter developed by Comma.ai [7] that provides access to almost all car sensors through the CAN bus and also enforces some safety constraints over output commands. However, when integrated with the CARLA driving simulator, OpenPilot does not utilize Panda software or hardware; thus Panda safety checks are inactive.
@@
Driver Reaction Simulator. To assess driver interventions, we develop a driver reaction simulator. The simulated driver is notified when any safety alerts are raised by the ADAS (e.g., FCW) or when the driver observes any abnormalities in the vehicle’s status or camera user interface (UI). These ACC abnormalities include hard braking, unexpected acceleration, the vehicle’s speed exceeding the cruising speed by more than 10%, or the mean perturbation value in the UI surpassing a noticeable threshold, with a default value of 15% representing an alert driver ($Patch.mean()>0.15$). We assume a very alert driver who can notice any anomalies that occur within a single control cycle (10ms). A predefined emergency response will be issued by the driver accordingly (see Table IV), taking effect 2.5 seconds later (average driver reaction time). To mimic the human driver’s braking behavior, we adopt a braking curve function from previous research [97].
TABLE IV: Driver simulator: activation conditions and reactions.
@@
Priority-based Control Command Dispatcher. With multiple safety mechanisms in place, there might be conflicts among the control commands issued by the OpenPilot ACC controller and those generated by the safety interventions. To resolve such conflicts, we design a command dispatcher to transmit output control commands to the CARLA actuators from various sources (e.g., ACC, AEB, simulated driver) based on their priorities, with high-priority commands overwriting low-priority ones (see Fig. 6 and Fig. 8). The simulated driver’s actions have a higher priority than regular ACC outputs, and control actions from the AEB have the highest priority. The driver’s actions will be executed 2.5s (average driver reaction time) after safety alerts (e.g., FCW) or noticing other ACC malfunctions (see Table IV). ACC commands will be blocked or disengaged when AEB or driver interventions are triggered.
Figure 7: AEBS.
Figure 8: Control command dispatcher.
V Evaluation in Simulated Environment
This section presents the evaluation of our proposed context-aware attack strategy (referred to as CA-Opt) using the enhanced simulation platform presented in Section IV.
V-A Methodology
We study the following research questions by comparing the effectiveness of CA-Opt attack to several baseline attack methods in causing safety hazards (Section V-B) and evading different safety interventions (Section V-C):
@@ V-A Methodology
Baselines. We design three baseline attack strategies to answer these questions (see Table V).
TABLE V: Overview of proposed and baseline attack strategies.
@@
To assess the effectiveness of our optimization-based adversarial patch method in strategically selecting the attack values, we compare it to two baselines: CA-Random, which introduces random perturbations to perception inputs, and CA-APGD, which uses a state-of-the-art gradient-based method, Auto-PGD [31], to determine the perturbation values. Since the original Auto-PGD is designed for misclassification, which does not work for attacking ACC, we change the goal function (to $\triangledown g(d)$, see Eq. 1) to maximize relative distance prediction. We also limit the iteration number to 5, the maximum number of control cycles (100Hz) within a perception cycle (20Hz)). Both baselines use the same context-aware method as the proposed CA-Opt attack strategy to choose the attack start time and duration. To evaluate our method’s efficiency in selecting the timing and duration of attacks, we design another baseline (Random-Opt) that selects a random start time uniformly distributed within [5, 40] seconds and a random attack duration uniformly distributed within [0.5, 2.5] seconds. To concentrate only on the effect of different start times and durations, Random-Opt shares the same adversarial patch generation method as CA-Opt. For a fair comparison, we confine perturbations to the detected bounding box (BBox) of the lead vehicle and update the BBox size and position with our proposed patch updating algorithm (Section 5).
Driving Scenarios. We model a 2016 Honda Civic, both with and without basic safety features, navigating curvy and straight sections of a highway using the "Town04_opt" map under clear weather and dry road conditions in CARLA. We simulate four high-risk driving scenarios designed based on the NHTSA’s pre-collision scenario topology report [67]. The Ego vehicle, traveling at 60 mph and positioned 75 meters away, encounters a lead vehicle exhibiting various behaviors: (SC1) Lead vehicle cruises at the speed of 35 mph; (SC2) Lead vehicle cruises at the speed of 50 mph; (SC3) Lead vehicle slows down from 50 mph to 35 mph; and (SC4) Lead vehicle accelerates from 35 mph to 50 mph.
@@
V-B Attack Success Rate in Causing Hazards
To assess the CA-Opt attack’s effectiveness in inducing safety hazards (RQ1), we conduct experiments in the closed-loop simulation platform without enabling the safety interventions. This setting is similar to previous works [76, 39].
We consider an attack successful if a collision event is observed (Ego vehicle collides with the lead vehicle) or the relative distance between the lead vehicle and the Ego vehicle is no larger than 0m. The success rate is reported across 1000 simulations, if not specified otherwise.
Figure 9: Top: Adversarial patch examples generated using our optimization method vs. random and APGD-based methods. Bottom: Success rate of CA-Opt and baseline attacks in absence of safety interventions.
Fig. 9 displays the success rates for each attack. The CA-Random attack leads to hazards in less than 1.2% of scenarios, with an overall success rate of 0.7%. This low rate suggests that randomly generated adversarial patches minimally impact the DNN model predictions and rarely cause hazards. Increasing the random perturbation values in the adversarial patch also does not significantly enhance the success rate. This is due to the DNN model in OpenPilot being primarily trained to detect any front objects, not to identify their specific class, thus showing greater resilience to adversarial attacks. Additionally, patches with larger perturbation values are more visible to the driver compared to those from the optimization-based method (Fig. 9-Top), potentially alerting the driver to prevent hazards.
In contrast, the proposed CA-Opt attack achieves a 100% success rate in all testing scenarios, surpassing CA-Random by 142.9 times. Although CA-APGD uses a similar goal function as CA-Opt, it does not cause hazards in 46.6% of simulations, mainly due to limiting its number of iterations to satisfy real-time constraints. On the other hand, by employing a dynamic patch updating algorithm, our proposed CA-Opt attack ensures the continuity of the optimization process across the perception cycles and enhances the attack effectiveness.
@@ V-B Attack Success Rate in Causing Hazards
V-C Attack Stealthiness with Safety Interventions
@@
V-C1 Stealthiness in Perception Input
To evade detection by safety mechanisms and human driver, the adversarial patch should stay as stealthy as possible. Basically, the smaller the value of the pixels’ perturbations, the stealthier the attack will be. Therefore, we tested our attack method with three different $\lambda$ values in Eq. 1. We use two sets of metrics to evaluate the stealthiness of the patch, including (i) the degree of the pixel perturbation measured using $L_{2}$ and $L_{\infty}$ distance [89] and (ii) the similarity between the original camera image and the perturbed image, calculated using RMSE and universal image quality index (UIQ) [86].
Table VI presents the results averaged over all the test scenarios and simulations. The CA-Opt attack achieves at least a 99.2% success rate under all three stealthiness levels and keeps the perturbation degree less than 0.015 ($L_{\infty}$) and 0.184 ($L_{2}$). The perturbed image with the adversarial patch has a similarity of UIQ = 0.993 (1 means identical) to the original image. We choose the $\lambda$ value to be $10^{-3}$ in our evaluations because of its stealthiness and high attack effectiveness. Examples of the generated adversarial patches (with $\lambda=10^{-3}$) are presented in Fig. 5 (see the zoomed-in area) and Fig. 9, which are almost invisible to human eyes.
TABLE VI: Attack success rate with different patch stealthiness levels $\lambda$ and pixel perturbation degrees.
@@ V-C1 Stealthiness in Perception Input
    $L_{2}$ and $L_{\infty}$ distances are the normalized perturbation values of the attack vector matrix in the range of [0,1].
@@
    Image similarity is evaluated by comparing the RMSE and UIQ between the original image and the perturbed image with the patch. Smaller RMSE and larger UIQ mean higher similarity.
V-C2 Evading Safety Interventions
For a more realistic evaluation of the effectiveness of different attack strategies, we rerun our experiments with different safety interventions (introduced in Section IV). A calibration of the safety features is performed before our evaluations to make sure the interventions are triggered correctly without any false positives.
@@ V-C2 Evading Safety Interventions
We test each attack method with different AEBS configurations: (i) FCW/AEB depends on an independent camera that is not compromised, (ii) AEB/FCW utilizes compromised camera inputs similar to the ACC (simulating stock ACC and AEBS that share a camera or independent ACC and AEBS cameras that are both compromised), or (iii) AEB/FCW is disabled. Driver intervention and ACC safety constraint checking (OpenPilot Panda checks) are considered for all three settings. Here, we do not test the Random-Opt attack due to its similarity to CA-Opt but with worse performance. We assess the efficacy of each attack method using metrics such as the attack success rate, safety intervention activation rate (indicating the percentage of simulations triggering safety interventions), and hazard prevention rate (the percentage of simulations where hazards occur without safety interventions).
@@
Table VII shows the experimental results of each attack method with different safety intervention configurations. We observe that, regardless of the interventions, the CA-Random and CA-APGD attacks fail to cause any hazards due to their low baseline success rates (see Fig. 9) and their noticeable perturbations that trigger the driver interventions in 23.8-27.4% and 100% of scenarios. These findings highlight the effectiveness of human drivers in preventing accidents and keeping autonomous driving safe.
@@
In contrast, with AEBS disabled (the last row of Table VII), the CA-Opt attack successfully evades driver intervention in all simulations. This is attributed to the small, virtually invisible value of the adversarial patch, resulting in an average attack success rate of 100%. We also conduct experiments that simulate higher driver sensitivity levels by decreasing the mean perturbation value threshold for activating driver intervention from the default value of 15% (see Section 6) to 10%, 5%, 2%, 1.5%, 1%, and 0.5%. As shown in Fig. 10, when perturbation thresholds are set to 0.5% and 1% (representing highly sensitive driver), the CA-Opt attack triggers driver interventions in all and 79.6% of the simulations, leading to attack success rates of 0% and 20.4%, respectively. However, with thresholds higher than 1.5%, our attack maintains a 100% success rate without triggering driver intervention. This finding underscores the robustness of the attack in evading driver detection across a range of driver sensitivities.
@@
TABLE VII: Performance of attacks with all the safety features and different AEBS settings.
@@
All & AEBS Compromised (Shared Camera) CA-Random 24.3% 0 100% (7/7)
@@
CA-Opt 14.6% 89.6% 10.4% (104/1000)
All & AEBS Disabled CA-Random 23.8% 0 100% (7/7)
CA-APGD 100% 0 100% (534/534)
CA-Opt 0 100% 0
Figure 10: Attack success rate and driver intervention activation rate with different driver sensitivity thresholds.
When the AEBS function is enabled and uses the same compromised camera inputs as ACC, CA-Opt attacks affect both the ACC and AEBS functionalities. So, the safety interventions are triggered in only 14.6% of simulations, leading to a high success rate of 89.6%. However, the CA-Opt attack encounters challenges when the AEBS relies on uncompromised camera data from an independent camera. In this scenario, the attack triggers safety interventions in all simulations. But it still maintains a success rate of 48.7% through a gradual (stealthy) change in the vehicle state (see Fig. 11) that delays AEBS activation and leaves insufficient time for hazard prevention.
@@
Figure 11: Context-Aware perception attacks vs. output attacks.
V-D Comparison to DNN Output and Control Output Attacks
The stealthy perturbations on the perception input can get propagated through the DNN model and ACC logic and lead to changes in the DNN output ( 2 in Fig. 1) and ACC control output 3. Although the attacker’s goal is to maximize errors in the DNN output and cause sudden accelerations on the ACC output, large deviations in vehicle states may be detected by the human driver or existing safety and defense mechanisms. To further evaluate the stealthiness of our proposed attack (CA-Opt), we compare the deviations resulting from the attack to those caused by stealthy attacks directly on DNN and control outputs. Note that such attacks are only possible under specific threat models (e.g., malware or wireless methods) in Table II.
Control Output Attacks. We first examine deviations in the autonomous vehicle states and control outputs resulting from the attack compared to two baseline output attacks, called MaxOut and StrategicOut. These attacks directly modify ACC output control commands, by setting them to a maximum allowed acceleration value (MaxOut) or a strategic value (StrategicOut) based on a method from prior research [97]. But they use the same context-aware method as CA-Opt for selecting the attack times and durations.

### v4

@@ 4. Safety Intervention Simulation
To evaluate the safety of DNN-based ACC systems under attacks, we enhance the default OpenPilot and CARLA simulation platform (see Section 2.3) to be more representative of real-world ADAS, by developing a safety intervention simulator and mechanisms for priority-based dispatching of control commands to CARLA and fusion of camera and radar data (see Appendix A). An overview of the simulation platform is shown in Fig. 6 (with the orange parts representing our new implementations) and presented next.
Figure 6. Simulation platform. [Code Available: https://github.com/gitguige/openpilot0.8.9]
@@
To fill the gap in considering safety interventions and address the challenge of ensuring the combination of human driver and vehicle safe (see Section 2.2), we implement and integrate three levels of safety interventions in the OpenPilot software (see Fig. 6), including ADAS safety features (e.g., AEB and FCW), basic car safety constraint checking on control commands, and driver interventions.
@@
AEBS (FCW and AEB) Simulator. To design and test the AEBS mechanisms in simulation, we thoroughly review the regulations and requirements concerning AEBS (schram_implementation_nodate, ) (http://data.europa.eu/eli/reg/2020/1597/oj, ) (noauthor_grva-12-50r1epdf_nodate, ) and adhere to UN Regulation No. 152 (http://data.europa.eu/eli/reg/2020/1597/oj, ). We adopt and implement a time-to-collision (TTC) based phase-controlled AEBS (alsuwian_autonomous_2022, ) in our platform.
@@
The AEBS processes inputs from LVD outputs (after sensor fusion), including relative distance ($RD$) and relative speed ($RS$), and current speed $V_{Ego}$ (see Fig. 1). The average driver reaction time ($T_{react}$) is set to the commonly used constant value of 2.5s (zhou2022strategic, ; sato2021dirty, ). Time thresholds, namely $ttc$ (time to collision), $t_{fcw}$ (FCW time), $t_{pb1}$ (1st phase partial brake time), $t_{pb2}$ (second phase partial brake time), and $t_{fb}$ (full brake time), are then calculated as follows:
(12) $ttc$ $=RD/RS;$
(13) $t_{fcw}$ $=T_{react}+V_{Ego}/4.5$
(14) $t_{pb1}$ $=V_{Ego}/2.8;t_{pb2}=V_{Ego}/5.8;t_{fb}=V_{Ego}/9.8$
@@
As shown int Fig. 8, when $ttc$ falls below $t_{fcw}$, $t_{pb1}$, $t_{pb1}$, and $t_{pb1}$, a corresponding action (warning or brake with 90%, 95%, 100% force) is executed. Applying the brake blocks other ADAS controls. See Appendix 9 for more details on AEBS design and testing.
@@
In reality, when OpenPilot is installed on a car, some car models lose the AEBS functionality (Carmodels-OpenPilot, ), while others retain it. Also, AEBS might rely on a separate ADAS camera (adas-camera-can, ), distinct from the OpenPilot camera, and a potential compromise of the AEBS camera data is possible. Thus, we consider three scenarios for AEBS interventions: (1) AEBS is enabled, and AEBS camera data is uncompromised; (2) AEBS is enabled, but AEBS camera data is compromised; and (3) AEBS is disabled (see Section 5.3.2 and Table 7).
@@
Safety Constraint Checker. The OpenPilot safety mechanisms are implemented in its control software and the Panda CAN interface. Panda is a universal OBD adapter developed by Comma.ai (Panda, ) that provides access to almost all car sensors through the CAN bus and also enforces some safety constraints over output commands. However, when integrated with the CARLA driving simulator, OpenPilot does not utilize Panda software or hardware; thus Panda safety checks are inactive.
@@
Table 4. Driver simulator: activation conditions and reactions.
@@
Driver Reaction Simulator. To assess driver interventions, we develop a driver reaction simulator. The simulated driver is notified when any safety alerts are raised by the ADAS (e.g., FCW) or when the driver observes any abnormalities in the vehicle’s status or camera user interface (UI). These ACC abnormalities include hard braking, unexpected acceleration, the vehicle’s speed exceeding the cruising speed by more than 10%, unsafe following distance with the lead vehicle (e.g., less than a vehicle length), or the mean perturbation value in the UI surpassing a noticeable threshold, with a default value of 15% representing an alert driver ($Patch.mean()>0.15$). We assume a very alert driver who can notice any anomalies that occur within a single control cycle (10ms). A predefined emergency response will be issued by the driver accordingly (see Table 4), taking effect 2.5 seconds later (average driver reaction time). To mimic the human driver’s braking behavior, we adopt a braking curve function from previous research (zhou2022strategic, ).
Figure 7. AEBS.
Figure 8. Control command dispatcher.
Priority-based Control Command Dispatcher. With multiple safety mechanisms in place, there might be conflicts among the control commands issued by the OpenPilot ACC controller and those generated by the safety interventions. To resolve such conflicts, we design a command dispatcher to transmit output control commands to the CARLA actuators from various sources (e.g., ACC, AEB, simulated driver) based on their priorities, with high-priority commands overwriting low-priority ones (see Fig. 6 and Fig. 8). The simulated driver’s actions have a higher priority than regular ACC outputs, and control actions from the AEB have the highest priority. The driver’s actions will be executed 2.5s (average driver reaction time) after safety alerts (e.g., FCW) or noticing other ACC malfunctions (see Table 4). ACC commands will be blocked or disengaged when AEB or driver interventions are triggered.
5. Evaluation in Simulated Environment
This section presents the evaluation of our proposed context-aware attack strategy (referred to as CA-Opt) using the enhanced simulation platform presented in Section 4.
5.1. Methodology
We study the following research questions by comparing the effectiveness of CA-Opt attack to several baseline attack methods in causing safety hazards (Section 5.2) and evading different safety interventions (Section 5.3):
@@ 5.1. Methodology
Baselines. We design three baseline attack strategies to answer these questions (see Table 5).
Table 5. Overview of proposed and baseline attack strategies.
@@
To assess the effectiveness of our optimization-based adversarial patch method in strategically selecting the attack values, we compare it to two baselines: CA-Random, which introduces random perturbations to perception inputs, and CA-APGD, which uses a state-of-the-art gradient-based method, Auto-PGD (croce2020reliable, ), to determine the perturbation values. Since the original Auto-PGD is designed for misclassification, which does not work for attacking ACC, we change the goal function (to $\triangledown g(d)$, see Eq. 1) to maximize relative distance prediction. We also limit the iteration number to 5, the maximum number of control cycles (100Hz) within a perception cycle (20Hz)). Both baselines use the same context-aware method as the proposed CA-Opt attack strategy to choose the attack start time and duration. To evaluate our method’s efficiency in selecting the timing and duration of attacks, we design another baseline (Random-Opt) that selects a random start time uniformly distributed within [5, 40] seconds and a random attack duration uniformly distributed within [0.5, 2.5] seconds. To concentrate only on the effect of different start times and durations, Random-Opt shares the same adversarial patch generation method as CA-Opt. For a fair comparison, we confine perturbations to the detected bounding box (BBox) of the LV and update BBox size and position with our proposed patch updating algorithm (Section 5).
Driving Scenarios. We model a 2016 Honda Civic, both with and without basic safety features, navigating curvy and straight sections of a highway using the "Town04_opt" map under clear weather and dry road conditions in CARLA. We simulate four high-risk driving scenarios designed based on the NHTSA’s pre-collision scenario topology report (najm_pre-crash_2007, ). The Ego vehicle, traveling at 60 mph and positioned 75 meters away, encounters a LV exhibiting various behaviors: (SC1) LV cruises at the speed of 35 mph; (SC2) LV cruises at the speed of 50 mph; (SC3) LV slows down from 50 mph to 35 mph; and (SC4) LV accelerates from 35 mph to 50 mph.
@@
5.2. Attack Success Rate in Causing Hazards
To assess the CA-Opt attack’s effectiveness in inducing safety hazards (RQ1), we conduct experiments in the closed-loop simulation platform without enabling the safety interventions. This setting is similar to previous works (sato2021dirty, ; guo2023adversarial, ).
We consider an attack successful if a collision event is observed (Ego vehicle collides with LV) or relative distance between LV and Ego vehicle is no larger than 0m. The success rate is reported across 1,000 simulations, if not specified otherwise.
Figure 9. Top: Adversarial patch examples generated using proposed method vs. random and APGD-based methods. Bottom: Success rate of CA-Opt and baseline attacks in absence of safety interventions.
Fig. 9 displays the success rates for each attack. The CA-Random attack leads to hazards in less than 1.2% of scenarios, with an overall success rate of 0.7%. This low rate suggests that randomly generated adversarial patches minimally impact the DNN model predictions and rarely cause hazards. Increasing the random perturbation values of the adversarial patch also does not significantly enhance the success rate. This is because OpenPilot’s DNN model is primarily trained to detect the front objects but not to identify their specific class, thus showing greater resilience to adversarial attacks. Additionally, patches with larger perturbation values are more visible to the driver than those from the optimization-based method (Fig. 9-Top), potentially alerting the driver to prevent hazards.
In contrast, the proposed CA-Opt attack achieves a 100% success rate in all testing scenarios, surpassing CA-Random by 142.9 times. Although CA-APGD uses a similar goal function as CA-Opt, it does not cause hazards in 46.6% of simulations, mainly due to limiting its number of iterations to satisfy real-time constraints. On the other hand, by employing a dynamic patch updating algorithm, the CA-Opt attack ensures the continuity of the optimization process across the perception cycles and enhances the attack’s effectiveness.
@@ 5.2. Attack Success Rate in Causing Hazards
5.3. Attack Stealthiness with Safety Interventions
@@
5.3.1. Stealthiness in Perception Input
To evade detection by safety mechanisms and human driver, the adversarial patch should stay as stealthy as possible. Basically, the smaller the value of the pixels’ perturbations, the stealthier the attack will be. Therefore, we tested our attack method with three different $\lambda$ values in Eq. 1. We use two sets of metrics to evaluate the stealthiness of the patch, including (i) the degree of the pixel perturbation measured using $L_{2}$ and $L_{\infty}$ distance (norm, ) and (ii) the similarity between the original camera image and the perturbed image, calculated using RMSE and universal image quality index (UIQ) (wang2002universal, ).
Table 6 presents the results averaged over all the test scenarios and simulations. The CA-Opt attack achieves at least a 99.2% success rate under all three stealthiness levels and keeps the perturbation degree less than 0.015 ($L_{\infty}$) and 0.184 ($L_{2}$). The perturbed image with the adversarial patch has a similarity of UIQ = 0.993 (1 means identical) to the original image. We choose the $\lambda$ value to be $10^{-3}$ in our evaluations because of its stealthiness and high attack effectiveness. Examples of the generated adversarial patches (with $\lambda=10^{-3}$) are presented in Fig. 5 (see the zoomed-in area) and Fig. 9, which are almost invisible to human eyes.
Table 6. Attack success rate with different patch stealthiness levels.
@@ 5.3.1. Stealthiness in Perception Input
    $L_{2}$ and $L_{\infty}$ distances are the normalized perturbation values of the attack vector matrix in the range of [0,1]. Image similarity is evaluated by comparing the RMSE and UIQ between the original image and the perturbed image with the patch. Smaller RMSE and larger UIQ mean higher similarity.
5.3.2. Evading Safety Interventions
@@
For a more realistic evaluation of the effectiveness of different attack strategies, we rerun our experiments with different safety interventions (introduced in Section 4). A calibration of the safety features is performed before the experiments to ensure the interventions are triggered correctly without any false positives. We test each attack method with different AEBS configurations: (i) AEB/FCW depends on an independent camera that is not compromised, (ii) AEB/FCW utilizes compromised camera inputs similar to the ACC (simulating stock ACC and AEBS that share a camera or independent ACC and AEBS cameras that are both compromised), or (iii) AEB/FCW is disabled. Driver intervention and ACC safety constraint checking (OpenPilot Panda checks) are considered for all three settings. Here, we do not test the Random-Opt attack due to its similarity to CA-Opt but with worse performance. We assess the efficacy of each attack method using metrics such as the attack success rate, safety intervention activation rate (indicating the percentage of simulations triggering safety interventions), and hazard prevention rate (the percentage of simulations where hazards occur without safety interventions).
@@
Table 7 shows the experimental results of each attack method with different safety intervention configurations. We observe that, regardless of the interventions, the CA-Random and CA-APGD attacks fail to cause any hazards due to their low baseline success rates (see Fig. 9) and their noticeable perturbations that trigger the driver interventions in 23.8-27.4% and 100% of scenarios. These findings highlight the effectiveness of human drivers in preventing accidents and keeping autonomous driving safe.
@@
In contrast, with AEBS disabled, the CA-Opt attack resulted in an average attack success rate of 82.6%. We also conduct experiments that simulate higher driver sensitivity levels by decreasing the mean perturbation value threshold for activating driver intervention from the default value of 15% (see Section 4) to 10%, 5%, 2%, 1.5%, 1%, and 0.5%. As shown in Fig. 10, when perturbation thresholds are set to 0.5% and 1% (representing highly sensitive driver), the adversarial patch triggers driver interventions in all and 79.6% of the simulations, leading to attack success rates of 0% and 20.4%, respectively. However, with thresholds higher than 1.5%, our attack maintains an 82.6% success rate. This finding underscores the robustness of the generated adversarial patch in evading driver detection across a range of driver sensitivities.
@@
Table 7. Performance of attacks with all the safety features and different AEBS settings.
@@
All & AEBS Disabled/ Compromised (Shared Camera) CA-Random 23.8%/ 24.3% 0 100% (7/7)
@@
CA-Opt 100% 82.6% 17.4% (174/1000)
Figure 10. Evaluation with different driver sensitivity thresholds.
When AEBS is enabled and uses the same compromised camera inputs as ACC, we observe a similar high success rate (82.6%) for the CA-Opt attacks that affect both the ACC and AEBS functionalities. However, the CA-Opt attack encounters challenges when the AEBS relies on uncompromised camera data from an independent camera. In this scenario, the attack triggers AEBS interventions in all simulations. But it still maintains a success rate of 48.7% through a gradual (stealthy) change in the vehicle state (see Fig. 11) that delays AEBS activation and leaves insufficient time for hazard prevention.
@@
5.4. Comparison to DNN Output and Control Output Attacks
The stealthy perturbations on the perception input can get propagated through the DNN model and ACC logic and lead to changes in the DNN output ( 2 in Fig. 1) and ACC control output 3. Although the attacker’s goal is to maximize errors in DNN output and cause sudden accelerations on the ACC output, large deviations in vehicle states may be detected by the human driver or existing safety and defense mechanisms. To further evaluate the stealthiness of our proposed attack (CA-Opt), we compare deviations resulting from the attack to those caused by stealthy attacks directly on DNN and control outputs. Note such attacks are only possible under specific threat models (e.g., malware or wireless methods) in Table 2.
Control Output Attacks. We first examine deviations in the autonomous vehicle states and control outputs resulting from the attack compared to two baseline output attacks, called MaxOut and StrategicOut. These attacks directly modify ACC output control commands, by setting them to a maximum allowed acceleration value (MaxOut) or a strategic value (StrategicOut) based on a method from prior research (zhou2022strategic, ). But they use the same context-aware method as CA-Opt for selecting attack times and durations.